how to shade the integrals

(b) 22 LIS (1, y) dy der Ji J3 f(x,y) dr dy
Answer #1

☺ se s }(my)dy de Here Y=3, y=y { x=1,7=2 yn - Y=4 - Save Posley - Y=3 q=1 Y=2s * S Pfungstedly 20 Hense a y=4 toy=5; x=0 to x=yal /x=y-1 aty=4 , x=3 - y25 9190 y=4 yr .53festady = sassurer + Sauf fryd